OpenStudy (anonymous):

The equation y = 55x + 85 describes the amount of money a plumber charges, where y is the total amount he charges and x is the number of hours he works. The table shows the total amount of money an electrician charges. How much does each charge per hour? Number of hours worked 1 2 4 7 Total amount charged 115 180 310 505 The plumber charges $_____per hour. The electrician charges $_____per hour.

OpenStudy (anonymous):

@Tootles143

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

I need help on this this confusing o.o

OpenStudy (anonymous):

All right i will try my best...

OpenStudy (anonymous):

@dan815 can you help... I am confused too...

OpenStudy (anonymous):

@Nnesha may you please help... I am confused too...

OpenStudy (anonymous):

This the last question and its confusing

Nnesha (nnesha):

find the slope of the table

OpenStudy (anonymous):

oh its 65

Nnesha (nnesha):

the equation they gave u is for plumber charges and table for electrician charges

Nnesha (nnesha):

that's right

OpenStudy (anonymous):

Sorry i wanst of help... :(

Nnesha (nnesha):

they are asking slope of equation and table how do i know they asking for slope bec there is key word "per" <-- slope

Nnesha (nnesha):

so The plumber charges $_____per hour. ????

OpenStudy (anonymous):

65

Nnesha (nnesha):

nope the equation they gave u is for plumber charges and table for electrician charges

Nnesha (nnesha):

The equation y = 55x + 85 describes the amount of money a plumber charge.

Nnesha (nnesha):

^^^rage of change slope ??

OpenStudy (anonymous):

? what

OpenStudy (anonymous):

The plumber charges $ 65per hour. The electrician charges $85 per hour.

Nnesha (nnesha):

85 where you got that ??

Nnesha (nnesha):

what is slope of the equation ??

OpenStudy (anonymous):

the equation

OpenStudy (anonymous):

i don't know how to solve the slope

Nnesha (nnesha):

yes

OpenStudy (anonymous):

the slope is 55

Nnesha (nnesha):

what?? what about previous post you just found slope there

OpenStudy (anonymous):

the other one was 65

Nnesha (nnesha):

equation y = mx +b here m is slope which is always with x

OpenStudy (anonymous):

The slope for this question is 55 y = mx +b y = 55x + 85

OpenStudy (anonymous):

equation*

Nnesha (nnesha):

yes that's right

Nnesha (nnesha):

what is slope of table ??

OpenStudy (anonymous):

65

Nnesha (nnesha):

yes so got the answer :)

OpenStudy (anonymous):

:) 65 goes on the bottom and 55 goes at the top?

OpenStudy (anonymous):

?you there

Nnesha (nnesha):

yes

OpenStudy (anonymous):

65 at the bottom and 55 at the top?

Nnesha (nnesha):

yes

Nnesha (nnesha):

The equation y = 55x + 85 describes the amount of money a plumber charges <-- they told u plumber charges he table shows the total amount of money an electrician charges.

OpenStudy (anonymous):

Plumber charges $85 on top of a $55 per hour fee thus y=55x+85 Electrician charges 115 after 1 hour and 180 after 2 hours the difference between 1 hour and 2 hours for the electricians additional charge is the difference between his 1 hour charge and 2 hour charge (115-180=65) yay per hour will be the 1 hour charge minus the additional fee (115-65=50) so the plumbers hourly rate is 55 while the electricians is 50. The moral of the story..be a plumber for the extra 5 an hour

OpenStudy (isuckatschool43):

Plumber: 55 Electrician: 65 confirmed 100%